OpenStudy (anonymous):

HARD PSAT QUESTION - "X is an integer and rational number. IF Y is the time it takes for object O to move 5 feet, then how far does object O move in 5 hours at the same speed?" Pls help - thanks! :) (By the way, "There is not enough information to solve this problem" was answer choice E - I don't remember the others.

OpenStudy (phi):

speed * time = distance speed * Y hours(?)= 5 ft speed = 5 ft/Y hours 5/Y * 5 hours = X (you don't say what x is , but I assume it's the unknown distance) 25= X*Y if X is an integer we can get an infinite number of answers. All it takes is for Y= 1/n where n is an integer.